1. Consideramarketwithtwoidenticalfirms.Themarketdemandis

?(?) = 26 − 2?,
where ? = ?1 + ?2 indicates total output and ? is the market price. Suppose that

both firms have the following identical total cost function:

??(??) = 2?? ,? = 1,2.

  1. Computethedifferencebetweenaverageandmarginalcostsforbothfirms.

  2. DerivetheCournotbest-responsefunctions.

  3. Solve for the Cournot-Nash equilibrium. Calculate the quantity, price and profit for each firm.

  4. Nowsupposethat,duetopoorinvestmentdecisions,thetotalcostfunctionof firm 1 changes to:

    ??(?1) = 2.5?1.

    The cost function of firm 2 is the same as before. Compute the Cournot-Nash equilibrium in this case and compare it with the equilibrium derived under “1c” above. What is the intuition behind your finding?

From the book Industrial Organization as I am currently studying for an exam
Answer #1

1) The market demand

P=26-2Q=26-2(q1+q2)

TC for firm 1=2q1, MC=2 (dTC/dq1), AC=2 (2q1/q1)

Similarly TC For firm 2=2q2, MC=2, AC=2

Therefore the difference between average and marginal cost for both firms is 0

TR for firm 1=Pq1 = q1(26-2(q1+q2))

MR for firm 1=dTR/dq1= 26-4q1-2q2

For profit maximization for firm 1:
MC=MR

1) 2=26-4q1-2q2

Thus best response for firm 1 when Firm 2 produces q2

is q1=(24-2q2)/4= 6-0.5q2

Similarly for firm 2:

TR for firm 2=Pq2 = q2(26-2(q1+q2))

MR for firm 1=dTR/dq2= 26-2q1-4q2

For profit maximization for firm 2:
MC=MR

2) 2=26-2q1-4q2

Thus best response for firm 2 when Firm 1 produces q1

is q2=(24-2q2)/4= 6-0.5q1

For both 1) and 2) To be true, we solve them simaltaneously to get

48=6q1+6q2 (From adding both equations)

and q1=q2(From subtracting one equation from another)

Therefore q1=q2=48/12=4

This is the cournot Nash equilibrium where both q1=q2=4, their best response to each others quantity will also be 6-0.5q=6-2=4 thus it is a Nash equililbrium

Now the cost function for firm 2 is same as before so the profit maximization condition MC=MR will remain same as above which is

2) 2=26-2q1-4q2

However the cost function for firm 1 is TC=2.5q1

MC=2.5

Therefore the profit maximizing condition for firm 1 is

2.5=26-4q1-2q2

For both to be true, we add and subtract them to find

47.5=6q1+6q2

and

0.5=2q2-2q1 or

q2=q1+0.25

So 47.5=12q1+1.5

q1=46/12=3.833 and q2=5.33

This is the new Cournot Nash equilibrium and in this case the intuition is that the Firm 1 produces less at an equilibrium because it has a higher Marginal cost and since MR decreases as it increases its own quantity.the MC=MR condition will occur sooner for first 1 and thus it will produce less than firm 2, and since firm 1 now produces less, firm 2 will have a higher value of q2 because q1 will be less in the profit maximizing condition for Firm 2.

Hope it's clear and helps. Do ask for any clarifications required.